use a linearization to estimate sin(pie+1/1000)
find the taylor polynomial of third degree of sin(x) centered at a=x
Answer #1

is given by :- к Linearization of function y = f(x) about a is, y = f(a) + f(a) (x-a). for sin (2 + 1 1 800) 1 ooo. f(x) = si